# A Unified Approach to Extract Intepretable Rules from Tree Ensembles via Integer Programming

### v1

@@ 5.2 Tabular classification
Similarly to the previous experiment, we compare the proposed method to the state-of-the-art about rule extraction for classification tasks of tabular data. As competing methods, we consider the previously mentioned approaches used for regression tasks. However, the currently available implementation of NodeHarvest cannot tackle classification problems, whereas SIRUS can only handle binary classification tasks. Therefore, we opted to exclude NodeHarvest from this experiment and select datasets with no more than different classes. In particular, the experimental plan involves running each method on a selected benchmark of 14 datasets from the OpenML public repository [60], following the approach of earlier works [39, 62]. To assess the effectiveness of methods in extracting very short and sparse rules resembling a decision tree, we extracted rules from a random forest of 500 trees of depth 2, and fixed the maximum amount of selected rule lists to 4. Similar to the previous experiment, our method ran without validation or tuning steps. For competing methods, we followed the parameter tuning procedures described by their authors. We divided each dataset into 75/25 train-test split. We ran each regression task with 30 different random seeds. We report the average and standard deviation of the out-of-sample accuracy in Table 4. Based on the results, we observe that competing methods achieved statistically equivalent classification performance, as shown by the critical distance diagram depicted in Figure 4. In particular, the proposed approach outperforms RuleFit. We find these results satisfactory. Arguably, interpretable methods are preferred when they perform similarly than black box methods [51]. The median value of computing times of the proposed method is minute, with Gurobi’s average runtime being of the total. Compared to competing methods, it is less efficient, as it requires more computational resources. However, our approach is more flexible than competing methods, as it can handle different data structures.
@@
| Dataset | Our method | RuleFit | SIRUS | Full model |
| adult | 0.81 0.02 | 0.76 0.01 | 0.76 0.02 | 0.75 0.00 |
| bank-marketing | 0.88 0.00 | 0.88 0.00 | 0.88 0.00 | 0.88 0.00 |
| banknote | 0.87 0.03 | 0.86 0.02 | 0.87 0.03 | 0.90 0.02 |
| blood-transfusion | 0.74 0.03 | 0.75 0.03 | 0.76 0.02 | 0.75 0.02 |
| diabetes | 0.69 0.03 | 0.69 0.05 | 0.70 0.04 | 0.74 0.02 |
| FICO | 0.67 0.04 | 0.67 0.04 | 0.71 0.01 | 0.72 0.03 |
| heart-c | 0.73 0.03 | 0.77 0.06 | 0.77 0.05 | 0.83 0.02 |
| ilpd | 0.70 0.04 | 0.71 0.02 | 0.72 0.03 | 0.72 0.03 |
| ionosphere | 0.89 0.05 | 0.80 0.06 | 0.88 0.04 | 0.92 0.03 |
| MagicTelescope | 0.75 0.02 | 0.73 0.02 | 0.73 0.01 | 0.75 0.02 |
| mushroom | 0.88 0.02 | 0.88 0.03 | 0.90 0.01 | 0.87 0.01 |
| musk | 0.87 0.02 | 0.85 0.01 | 0.93 0.04 | 0.85 0.03 |
| tic-tac-toe | 0.68 0.02 | 0.66 0.02 | 0.67 0.03 | 0.68 0.01 |
| wdbc | 0.90 0.02 | 0.90 0.03 | 0.91 0.02 | 0.94 0.02 |
| Average | 0.79 0.03 | 0.78 0.03 | 0.80 0.03 | 0.81 0.02 |
| Rank () | 2.71 | 3.21 | 2.14 | 1.93 |
@@ 5.3 Time series classification
The aim of this experiment is to assess the effectiveness of the proposed method in handling tree ensembles specifically designed for time series data. We chose the generalized Random Shapelet Forest (gRSF) [38] as the tree ensemble method for this purpose. To the best of our knowledge, gRSF represents the only ensemble method for time series data that provides interpretable features (shapelets), and our approach is the only extraction method working on temporal tree ensembles. The experimental plan consists of running gRSF and our method on a selected benchmark of 25 time series datasets from the UCR public repository [24], following the methodology of Karlsson et al. [38]. We remark that the UCR repository provides fixed train-test splits for each dataset, obviating the need for additional splitting procedures. Consequently, we conducted each classification task using only 10 different random seeds. In this experimental evaluation, we regard gRSF as the Full model and apply the proposed method to extract lists of rules from it. The gRSF method was trained with the tuned hyperparameters provided by its authors, with the exception of fixing the tree depth to 3 to maintain consistency with the previous experiments. Therefore, we selected time series datasets with no more than different classes. To assess the effectiveness of the training procedure discussed in Section 4.3, we train our method by fixing parameter and by running the procedure with 5-fold cross-validation. The average and standard deviation of the out-of-sample accuracy are reported in Table 6. From the obtained results, the proposed method with the training procedure achieves statistically equivalent performance to gRSF (Wilcoxon’s test -value ), with an average of 15 extracted rule lists. Our method with a fixed obtains lower accuracy values but runs faster. The median computation time with the training procedure is minutes, whereas with fixed , it is minutes. On average, the runtime of the Gurobi IP solver is negligible. Remarkably, with given hyperparameters, gRSF is significantly faster, running in less than 20 seconds per task. Nevertheless, the computational overhead of our method is justified by its interpretability and condensed predictive power.
| Dataset | Train | Test | Length | Classes |
| size | size | |||
@@
| Dataset | Our method () | Our method (w/ tuning) | gRSF (Full model) |
| Beef | 0.55 0.09 | 0.68 0.10 | 0.65 0.05 |
| CBF | 0.85 0.07 | 0.93 0.06 | 0.98 0.00 |
| CinCECGTorso | 0.69 0.06 | 0.83 0.06 | 0.79 0.00 |
| Coffee | 0.87 0.08 | 0.96 0.04 | 0.93 0.01 |
| DiatomSizeReduction | 0.83 0.10 | 0.96 0.04 | 0.98 0.01 |
| ECG200 | 0.82 0.02 | 0.92 0.03 | 0.84 0.02 |
| ECGFiveDays | 0.82 0.12 | 0.90 0.05 | 0.77 0.01 |
| FaceFour | 0.82 0.08 | 0.95 0.02 | 0.96 0.00 |
| GunPoint | 0.83 0.07 | 0.92 0.03 | 0.95 0.01 |
| Haptics | 0.38 0.03 | 0.61 0.12 | 0.40 0.02 |
| InlineSkate | 0.32 0.03 | 0.74 0.08 | 0.34 0.02 |
| ItalyPowerDemand | 0.95 0.01 | 0.93 0.02 | 0.95 0.01 |
| MoteStrain | 0.76 0.09 | 0.85 0.06 | 0.91 0.00 |
| OliveOil | 0.79 0.07 | 0.89 0.04 | 0.87 0.02 |
| Sony.1 | 0.87 0.10 | 0.96 0.03 | 0.89 0.01 |
| Sony.2 | 0.86 0.05 | 0.88 0.04 | 0.92 0.01 |
| StarLightCurves | 0.93 0.03 | 0.97 0.05 | 0.91 0.01 |
| Symbols | 0.61 0.11 | 0.76 0.06 | 0.92 0.01 |
| Trace | 0.96 0.09 | 0.96 0.05 | 0.82 0.00 |
| TwoLeadECG | 0.87 0.05 | 0.91 0.03 | 0.96 0.00 |
| TwoPatterns | 0.62 0.11 | 0.77 0.20 | 0.86 0.00 |
| UWaveGestureX | 0.62 0.04 | 0.61 0.13 | 0.73 0.02 |
| UWaveGestureY | 0.53 0.04 | 0.59 0.11 | 0.61 0.01 |
| UWaveGestureZ | 0.58 0.04 | 0.77 0.22 | 0.62 0.00 |
| Wafer | 0.94 0.02 | 0.98 0.03 | 0.90 0.00 |
| Average | 0.75 0.06 | 0.85 0.07 | 0.82 0.01 |
6 Conclusion and future work
In this paper, we introduce a novel method to extract interpretable rule lists from tree ensembles. Our approach involves solving a set partitioning problem, separating the computation of loss and regularization functions from the Integer Programming model. Leveraging the flexibility of Mathematical Programming, our method can incorporate multiple non-linear functions and it can handle both classification and regression tasks for tabular and time series data. We present extensive and statistically significant computational results for the classification and regression of tabular data, showing that the proposed method achieves comparable performance to other competing rule-extraction methods for tree ensembles. Notably, our approach works as a parameter-free method under common benchmark settings. In cases where prior task-specific knowledge is absent and the number of rule lists cannot be inferred beforehand, we propose an automated training procedure to validate our method effectively. Additionally, we provide empirical evidence that the proposed method can effectively extract rules from tree ensemble methods for the classification of temporal data. To the best of our knowledge, this is the first rule extraction method capable of handling complex data such as time series. Experiments also show that our method is computationally demanding and less efficient than competing approaches. Nevertheless, this behavior is almost completely ascribable to the preprocessing steps, empirically proving the efficiency of the proposed IP model. Future research may focus on improving computational efficiency and extending the introduced formalism to handle tree ensembles for more complex tasks, such as image [11, 45, 64] or text classification [12, 56].
@@ Acknowledgments
This work was supported by Fedegari Autoclavi S.p.A. The authors thank Prof. Stefano Gualandi and Prof. Isak Samsten for their valuable support.

### v3

@@ 5.2 Tabular classification
Similarly to the previous experiment, we compare the proposed method to the state-of-the-art about rule extraction for classification tasks of tabular data. As competing methods, we consider the previously mentioned approaches used for regression tasks. However, NodeHarvest and SIRUS can only handle binary classification tasks. Therefore, we select datasets with no more than different classes. In particular, the experimental plan involves running each method on a selected benchmark of 14 datasets from the OpenML public repository [82], following the approach of earlier works [56, 87]. To assess the effectiveness of methods in extracting very short and sparse rules that resemble a decision tree, we extract rules from a random forest of 500 trees of depth 2, and fix the maximum amount of selected rules to 4. Similar to the previous experiment, this quantity coincides with parameter , making the proposed validation procedure unnecessary. For competing methods, we fix the number of extracted rules to 4 and the maximal interaction order to 2. NodeHarvest does not need further parameter tuning. For SIRUS, we tune the frequency threshold through cross-validation, and we set the parameter to 10. For RuleFit, we tune the regularization parameter for the logistic model through cross-validation. We divide each dataset into a 75/25 train-test split. We run each classification task with 30 different random seeds, reporting the average and standard deviation of the out-of-sample accuracy in Table 5. From the critical difference diagram comparing ranks in Figure 6, we observe that all methods achieve statistically equivalent classification performance, with the exception of NodeHarvest, which performs poorly when running with a very limited number of nodes. The median value of computing times of the proposed method is minute, with Gurobi’s average runtime being of the total. Compared to competing methods, it is less efficient, as it requires more computational resources. However, our approach is more flexible than competing methods, as it can handle different loss functions and multi-class classification tasks.
@@
Dataset | Our method | RuleFit | SIRUS | NH |
adult | 0.81 0.01 | 0.77 0.02 | 0.76 0.02 | 0.75 0.00 |
bank-mark. | 0.89 0.00 | 0.89 0.01 | 0.88 0.00 | 0.88 0.00 |
banknote | 0.91 0.01 | 0.85 0.08 | 0.87 0.03 | 0.56 0.02 |
blood-transf. | 0.74 0.02 | 0.76 0.02 | 0.76 0.02 | 0.75 0.03 |
diabetes | 0.70 0.03 | 0.70 0.05 | 0.70 0.04 | 0.65 0.02 |
FICO | 0.69 0.01 | 0.67 0.03 | 0.71 0.01 | 0.52 0.01 |
heart-c | 0.73 0.04 | 0.75 0.06 | 0.77 0.05 | 0.54 0.07 |
ilpd | 0.70 0.03 | 0.72 0.03 | 0.72 0.03 | 0.72 0.03 |
ionosphere | 0.89 0.02 | 0.83 0.07 | 0.88 0.04 | 0.70 0.09 |
MagicTeles. | 0.77 0.02 | 0.74 0.03 | 0.73 0.01 | 0.65 0.01 |
mushroom | 0.91 0.02 | 0.88 0.07 | 0.90 0.01 | 0.62 0.01 |
musk | 0.86 0.01 | 0.87 0.02 | 0.93 0.04 | 0.85 0.01 |
tic-tac-toe | 0.68 0.02 | 0.67 0.05 | 0.67 0.03 | 0.65 0.02 |
wdbc | 0.95 0.03 | 0.92 0.03 | 0.91 0.02 | 0.63 0.03 |
Average | 0.80 0.02 | 0.79 0.04 | 0.80 0.03 | 0.68 0.03 |
Rank () | 1.89 | 2.25 | 2.11 | 3.75 |
5.2.1 Comparison in terms of fidelity
To assess internal fidelity, we compute the same measures as in the previous experiment. For external fidelity, we calculate the disagreement as the fraction of misaligned classifications between each competing method and the random forest. We report in Table 6 the average of each fidelity measure computed over the benchmark datasets, for each rule extraction method. For the sake of completion, we also report the average number of extracted rules. From obtained results, it turns out that the proposed method excels in representativeness, while obtaining the second best F1-score. Instead, SIRUS achieves the lowest disagreement value and the best F1-score, but performs poorly in terms of representativeness. This behavior is ascribable to its tendency to extract rules with a limited number of logical conditions, as shown in B.
Measure | Our method | RuleFit | SIRUS | NH |
Represented trees | 0.355 | 0.178 | 0.064 | 0.112 |
Represented paths | 0.021 | 0.008 | 0.006 | 0.005 |
F1-score | 0.546 | 0.259 | 0.571 | 0.152 |
Disagreement | 0.107 | 0.097 | 0.054 | 0.179 |
Extracted rules | 3.98 | 3.24 | 3.61 | 2.48 |
@@ 5.3 Time series classification
The aim of this experiment is to assess the effectiveness of the proposed method in handling tree ensembles specifically designed for time series. We choose the generalized Random Shapelet Forest (gRSF) [52] as the tree ensemble model for this purpose, providing interpretable features (shapelets). Currently, our approach is the only rule extraction method whose implementation works on temporal data. The experimental plan consists of running gRSF on a selected benchmark of 27 time series datasets from the UCR public repository [31] following the methodology of Karlsson et al. [52], and applying the proposed method to extract an optimal list of rules from it. We remark that the UCR repository provides fixed train-test splits for each dataset, avoiding the need for additional splitting strategies. For this reason, we conduct each classification task using only 10 different random seeds. The gRSF is trained using the tuned hyperparameters provided by its authors, which control the size and number of sampled shapelets. To maintain consistency with the previous experiments, we fix the maximal depth to 3. Therefore, we select time series datasets with no more than different classes. As the number of rules to extract is not fixed, we train our method by running the validation procedure with exact bounds discussed in Section 4.3, and we apply 5-fold cross-validation. The average and standard deviation of the out-of-sample accuracy are reported in Table 8. On average, results show that the gRSF (Full model) outperforms the proposed method, while the latter manages to extract an average of 6 interpretable rules. The median computation time with the training procedure is minutes. On average, the runtime of the Gurobi IP solver is negligible. Remarkably, with given hyperparameters, gRSF is significantly faster, running in less than 30 seconds per task. Nevertheless, the computational burden of our method is justified by its interpretability and condensed predictive power. In B, we show an example of a shapelet-based list of rules extracted from gRSF.
Dataset | Train size | Test size | Length | Classes |
@@
Dataset | Our method | gRSF (Full model) |
Beef | 0.45 0.10 | 0.70 0.02 |
CBF | 0.83 0.06 | 0.99 0.00 |
ChlorineConcentration | 0.56 0.01 | 0.58 0.00 |
CinCECGTorso | 0.62 0.05 | 0.79 0.00 |
Coffee | 0.87 0.06 | 0.90 0.02 |
DiatomSizeReduction | 0.83 0.05 | 0.96 0.01 |
ECG200 | 0.81 0.05 | 0.84 0.01 |
ECGFiveDays | 0.97 0.03 | 1.00 0.00 |
FaceFour | 0.74 0.07 | 0.98 0.00 |
GunPoint | 0.88 0.06 | 0.99 0.00 |
Haptics | 0.39 0.03 | 0.43 0.02 |
ItalyPowerDemand | 0.93 0.01 | 0.94 0.00 |
Lightning2 | 0.77 0.03 | 0.79 0.01 |
Lightning7 | 0.55 0.05 | 0.66 0.01 |
MoteStrain | 0.73 0.03 | 0.87 0.01 |
OliveOil | 0.69 0.05 | 0.86 0.01 |
SonyAIBORobotSurface1 | 0.91 0.05 | 0.96 0.00 |
SonyAIBORobotSurface2 | 0.86 0.04 | 0.94 0.00 |
Symbols | 0.61 0.10 | 0.92 0.01 |
SyntheticControl | 0.93 0.02 | 0.99 0.00 |
Trace | 0.95 0.02 | 0.99 0.00 |
TwoLeadECG | 0.91 0.03 | 0.95 0.00 |
TwoPatterns | 0.62 0.04 | 0.86 0.01 |
UWaveGestureX | 0.63 0.03 | 0.70 0.00 |
UWaveGestureY | 0.53 0.03 | 0.62 0.00 |
UWaveGestureZ | 0.78 0.02 | 0.64 0.01 |
Wafer | 0.96 0.01 | 0.97 0.00 |
Average | 0.74 0.04 | 0.84 0.01 |
6 Conclusions and future work
In this paper, we introduce a novel method to extract interpretable rules from tree ensembles. Our approach involves solving a set partitioning problem, separating the computation of loss and regularization functions from the Integer Programming model. Leveraging the flexibility of Mathematical Programming, our method can incorporate various non-linear functions, and it can handle both multi-class classification and regression tasks for tabular and temporal data. We present extensive and statistically significant computational results for the classification and regression of tabular data, showing that the proposed method achieves comparable performance to other competing rule extraction methods for tree ensembles. We also conduct a comparison in terms of explanation fidelity, introducing two novel measures of internal fidelity for surrogate models from tree ensembles. The results demonstrate that the proposed approach outperforms others in internal fidelity, while maintaining competitive external fidelity. Additionally, we provide empirical evidence that the proposed method can effectively extract interpretable rules from tree ensembles designed for the multi-class classification of temporal data. Experiments also show that our method is computationally demanding and less efficient than competing approaches. Nevertheless, this behavior is almost completely ascribable to the preprocessing steps, empirically proving the efficiency of the proposed IP model. Future research may focus on improving the preprocessing phase and extending the implementation of the proposed method to handle tree ensembles for more complex tasks, such as image [14, 63, 89] or text classification [15, 77].
@@ Acknowledgments
This work was supported by Fedegari Autoclavi S.p.A. The authors thank Prof. Stefano Gualandi and Prof. Isak Samsten for their valuable support. The authors express their gratitude to all the researchers who generously provided the datasets used in this work.
Appendix A Comparison of rule extraction methods
We note that competing methods can generate significantly different types of rules. RuleFit is the only method that produces rules similar in form to those generated by our approach. In contrast, SIRUS generates rules in an “if-then-else” format, where each rule could be interpreted as two separate rules according to our definition, due to the inclusion of the “else” part. Finally, NodeHarvest generates rules in the form of weighted trees, where each node contributes to the prediction and is weighted depending on the tree structure. In the case of NodeHarvest, branch nodes of weighted trees can be viewed as rules, as they may become leaf nodes after the pruning step. As a result, ensuring a fair comparison across methods is not trivial. Nonetheless, each of the methods considered allows for direct control over the upper limit of rules to be extracted. Therefore, in our experimental evaluation we choose to maintain the same fixed number of rules according to the corresponding definitions used by the competing methods, as specified by their authors.
Appendix B Interpretable output
We present examples of the interpretable rules extracted by different methods. The target datasets selected for this qualitative analysis are those from the benchmark where all methods performed similarly. For simplicity, we consider RuleFit and SIRUS as competing methods, since their lists of rules are easier to represent.
Tecator
For this tabular dataset, we trained a random forest with maximal depth 2 and we limited the number of extracted rules up to 4. We do not apply any train-test split, as we are not evaluating out-of-sample performance. For completeness, we report the prediction values of SIRUS rules obtained by the “else” clause. In Table 9, we present the lists of rules obtained from our method, RuleFit, and SIRUS, along with their corresponding weights and prediction values. Notably, the rules extracted by our method are the only ones that construct a complete decision tree, as depicted in Figure 7. Instead, RuleFit tends to extract logical conditions with slightly different thresholds, and SIRUS tends to produce shorter rules, due to the inclusion of “else” conditions.
Our method | |||
Rule | Weight | Prediction | |
- | 2.12 | ||
- | 0.84 | ||
- | -0.13 | ||
- | -0.80 | ||
RuleFit | |||
Rule | Weight | Prediction | |
1.78 | 2.21 | ||
0.54 | 0.87 | ||
-0.45 | -0.54 | ||
-0.67 | -0.80 | ||
SIRUS | |||
Rule | Weight | Prediction | Prediction (else) |
0.57 | 1.32 | -0.57 | |
0.44 | 2.16 | -0.24 | |
0.27 | 1.19 | -0.50 |
Houses
For this tabular dataset, we trained a random forest with maximal depth 3 and we limited the number of extracted rules up to 15. We do not apply any train-test split, as we are not evaluating out-of-sample performance. For completeness, we report the prediction values of SIRUS rules obtained by the “else” clause. In Table 10, we present the lists of rules obtained from our method, RuleFit, and SIRUS, along with their corresponding weights and prediction values. Again, the rules extracted by our method are the only ones that form a complete decision tree, as shown in Figure 8. Instead, RuleFit tends to extract shorter rules containing logical conditions with slightly different thresholds, while SIRUS tends to produce more rules made of only one condition.
Our method | |||
Rule | Weight | Prediction | |
- | -0.42 | ||
- | -0.79 | ||
- | -0.08 | ||
- | 0.45 | ||
- | 0.64 | ||
- | 1.35 | ||
- | 1.45 | ||
- | 2.17 | ||
RuleFit | |||
Rule | Weight | Prediction | |
-0.33 | -0.67 | ||
0.38 | 1.06 | ||
-0.36 | -0.60 | ||
1.02 | 1.86 | ||
-0.05 | -0.28 | ||
-0.34 | -0.24 | ||
-0.04 | -0.27 | ||
SIRUS | |||
Rule | Weight | Prediction | Prediction (else) |
0.41 | -0.28 | 1.11 | |
0.45 | 0.08 | -0.75 | |
0.17 | 0.33 | -0.08 | |
0.12 | -0.04 | 0.35 | |
0.53 | -0.60 | 0.40 | |
0.11 | -0.13 | 0.20 | |
0.29 | -0.33 | 0.79 | |
0.12 | -0.20 | 0.13 | |
0.11 | -0.16 | 0.16 | |
0.07 | 0.10 | -0.06 | |
0.21 | -0.34 | 0.63 |
ECGFiveDays
This temporal dataset contains ECG signals recorded from a 67-year-old male. The two classes correspond to the dates 12/11/1990 and 17/11/1990 [31]. For this dataset, we first trained the gRSF over the training set by using the provided hyperparameters [52] and setting the maximal depth to 3. Since no competing methods currently address time series classification, we applied our rule extraction method with the parameter fixed at 2. The resulting list of rules is shown in Figure 9, presented as a shapelet-based decision tree with unitary depth. The first leaf node corresponds to the date 17/11/1990, while the second leaf node corresponds to 12/11/1990. We also present two examples of time series from different classes, overlaying the single shapelet extracted from the gRSF and reporting the distance values computed using the subsequence distance measure (1). Notably, the temporal positions of the two peaks near are sufficient for accurately categorizing the signals. The extracted rules effectively capture and explain the predictions of the shapelet-based tree ensemble, obtaining a test accuracy of .